OpenStudy (anonymous):

The total cost in dollars to produce q units of a product is C(q). Fixed costs are $13000 . The marginal cost is C'(q)=0.006q^2-q+50. Estimate C(170) , the total cost to produce 170 units.Find the value of C'(170). Then combine these to find C(171). I found the value of C'(170) but am unsure of how to get the original equation to help solve the rest of the problem.

OpenStudy (anonymous):

I am assuming C'(q) does not mean derivative

OpenStudy (anonymous):

it does.

OpenStudy (anonymous):

\[(\int\limits_0^{170} \left(0.006 q^2-q+50\right) \, dq\]

OpenStudy (anonymous):

so I don't need to find the original function?

OpenStudy (anonymous):

you are not given C[q]

OpenStudy (anonymous):

but I don't need to solve for it?

OpenStudy (anonymous):

you need to integrate then add it to fix costs of 13,000

OpenStudy (anonymous):

I got 3876 from integration

OpenStudy (anonymous):

i got 3876 as well. then i find the derivative.and for 171, add to 13000?

OpenStudy (anonymous):

integrate up to 171 to get c(171)

OpenStudy (anonymous):

after i integrate, add 13000?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i might have integrated wrong then because I got 9070.078 after adding 13000.

OpenStudy (anonymous):

are you in diff eq class

OpenStudy (anonymous):

no im in calc 1.

OpenStudy (anonymous):

ok, good

OpenStudy (anonymous):

i did the same steps for the next problem but i keep getting the first and last part worng but the middle correct.

OpenStudy (anonymous):

wrong*

OpenStudy (anonymous):

C[170]=16876 is this right?

OpenStudy (anonymous):

Then the question says add C'(170) which I found to be 53.4 totalling 16929.40. Is that correct according to your answers?

OpenStudy (anonymous):

yes that it correct. how'd you get that value exactly?

OpenStudy (anonymous):

\[\int\limits_0^{170} \left(0.006 q^2-q+50\right) \, dq\]=3876 3876 + 13000=16876 add C'[170]=53.4 16876+53.4=16929.40

OpenStudy (anonymous):

I plugged 170 into the C'(q) equation. That gave me 53.4(well 53.40 in dollars that is). If you add that amount to C(170) it basically increases you to how much it would cost to produce 171 units. Now if you plugged C'(171) in and added that amount to 16929.40 it would give the amount of dollars to produce 172 units. See what we're doing? It's like voodoo! LOL

OpenStudy (anonymous):

oh that's the part I missed! Thanks so much for helping me out& posting really, I appreciate it a lot!

OpenStudy (anonymous):

Welcome of course!

OpenStudy (anonymous):

:)